OpenStudy (anonymous):

Which line is a graph of the equation: 2x + 5y = 10? Number graph ranging from negative five to five on the x axis and negative six to two on the y axis. Four lines are drawn in blue on the graph. Lines a and c have a positive slope and are parallel to each other. Lines b and d have a negative slope and are parallel to each other. Lines a and b intersect at (zero, two), and lines c and d intersect at (zero, negative two). A. line a B. line b C. line c D. line d

OpenStudy (photon336):

Well @gemelita1 i'm not sure about what you've written because i'd need to see a figure. But what I can do is re-write the equation in the form: \[Y = m(x)+b\] first we isolate y by subtracting 2x from both sides. \[2x+5y = 10\] 5y+2x = 10 -2x -2x 5y = 10-2x 5 5 5 Then we divide both sides by 5 y = 2-(2/5x)

OpenStudy (anonymous):

thanks :)